OpenStudy (anonymous):

Solve the following system using the substitution method. x + 3y = –5 5x + 6y = –19 a. negative three, negative two–thirds b. negative three, two–thirds c. three, negative two–thirds d. three, two–thirds

OpenStudy (anonymous):

a

OpenStudy (anonymous):

are you sure?

jimthompson5910 (jim_thompson5910):

Start with x + 3y = –5 and isolate x to get x + 3y = –5 x = –5-3y

jimthompson5910 (jim_thompson5910):

Then move onto 5x + 6y = –19 and replace 'x' with -5-3y since x = -5-3y to get 5x + 6y = –19 5(-5-3y) + 6y = –19

jimthompson5910 (jim_thompson5910):

Solve that for y. Then use that solution of y to find the solution for x.

OpenStudy (anonymous):

-25 - 15y + 6y = -19 -25 - 9y = -19 +25 + 25 -9 = 6

OpenStudy (anonymous):

@jim_thompson5910 6 2 -- = --- -9 -3

jimthompson5910 (jim_thompson5910):

yes, y = -2/3

jimthompson5910 (jim_thompson5910):

now find x

OpenStudy (anonymous):

im not good with fractions inside of equations

jimthompson5910 (jim_thompson5910):

x + 3y = –5 x + 3(-2/3) = –5 x + (3/1)(-2/3) = –5 x + (3(-2))/(1*3) = –5 x + (-6/3) = -5 x - 2 = -5 I'll let you finish

OpenStudy (anonymous):

x - 2 = -5 + 2 + 2 x = -3

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

THANK YOU!

jimthompson5910 (jim_thompson5910):

np